Team Avengers were fighting a group of villains in which they have superpowers to stop  a bullet with their hands. From your observation, what is the possible force that the villains  used to stop the firing bullet? Assume that the bullet mass is 15.0 g and it is travelling at  130 m/s as it penetrates the villains’ hands for about 2 cm. Bear in mind that the villains’  hand is quite flexible where the force deviated at angle of 10°. Give your answer to correct  two significant figures.


1
Expert's answer
2021-04-19T17:06:15-0400

First, find the magnitude of the force that villains' hands will meet:


"f=F\cos\theta."

Now apply Newton's second law:


"f=ma=m\frac{v^2}{2d}=F\cos\theta,\\\space\\
F=\frac{mv^2}{2d\cos\theta}=6435\text{ N}."